Question 4 (force vectors) Two people are pulling on a crate through two ropes. Person A is on the left pulling toward the left with a force of 50 N at 45° and person B is on the right pulling toward the right at an angle of 30%. Here is what the situation looks like from above. The resulting acceleration of the crate is at 0°. acceleration direction person B F@30% person A , 50 N 45o (30% 45°! crate a. What is the magnitude of the force from person B on the crate? b. What is the angle of the force from person A on the crate? If a person C comes in and starts to push the crate to the left from the right with 10 N of force, acceleration direction person B F@30% person A. 50 N 45; 30% 45° person C 10 N crate c. Can person B adjust the value of F to keep the crate accelerating in the same direction? If so, how much force must F be?
Answer #1

A for the force to be along 0 deg the forces in the horizontal direction should be canceled by each other so the horizontal component of the force from person A is 50sin45 and the horizontal component of the person B is F sin 30 then Fsin30 =50 sin 45 then force F = 50 sin45/sin30 = 70.71N

B from the question the angle is 45 degrees with the crate

C again applying equilibrium equation along the horizontal axis 50sin45 + 10 = F sin 30 ==> 35.355+10 = f*0.5

F = 45.355 * 2 = 90.71N and yes the person B can adjust the force to keep the net acceleration of the crate along 0 degrees axis